a) A laser produces a beam of light whose average power is 4W. How much energy is stored in a 3m length of the beam? Use c=3*10^8

b) If the area of the beam in part a) is .001m^2, what is the average intensity of the beam?

c) If unpolarized light had an average intensity equal to that found in part b), and it were passed first through a polarizer, then through an analyzer whose axis of transmission is tilted 45 degrees with respect to the axis of transmission of the polarizer, what is the intensity of the light upon leaving the analyszer?

d) For an electromagnetic wave of average intensity S, what is the peak magnetic field of the wave?

e) If the frequency of a laser were to double, what would happen to the wavelength of the laser light?
Answer #1

Enery per unit length = Power/c
   = 4/3*10^8

= 1.333*10^-8 J/m

a) enrgy in 3m, E = 1.333*10^-8*3 = 4*10^-8 J

b) Intensity = power/Area

= 4/0.001

= 4000 W/m^2

c) Io = 4000/2 = 2000 w/m^2

I = Io*cos^(45)

= 2000*0.5

= 1000 w/m^2

d)

e) c = lamda*f

c is always constant. so, if frequnecy is doubled wavelnegth becomes half.